Fermi National Laboratory
 

Accelerator Update for May 18, 2007 - May 21, 2007

 

Friday May 18

The day shift began with Operations monitoring the Tevatron (TeV) store 5435 (48.73E30) and the Recycler stash (176.2E10), with the Antiproton Source (Pbar) stacking (34.1E10), with the Linac, Booster, and Main Injector (MI) operating normally, with NuMI and MiniBooNE experiments taking beam, and with D-Zero (D0) and the Collider Detector at Fermilab (CDF) using the colliding beam.

At 9 AM, the MiniBooNE beam permit tripped off due to bad beam positions. Operators lowered the experiment’s intensity and tuned.

At 10:10 AM, the NuMI beam permit tripped off due to a power supply trip (VP121). Everything reset.

At 10:43 AM, a Booster vacuum pump tripped off and caused two RF stations (BRF15 & 16) to trip off. Everything reset.

At 12:59 PM, a Linac RF station (LRF3) tripped off. It reset.

At 1:23 PM, a Pbar Accumulator RF station (ARF1) tripped off. It reset.

Operations terminated store 5435 at 3 PM.

Operations put the TeV into a wet squeeze at 3:56 PM for a proton only study.

Operations reported at 3:57 PM that the Recycler had stashed antiprotons three times during the day shift.

TeV experts completed their study at 4:24 PM and Operations put the TeV into shot setup.

At 5:55 PM, Operations reported that the Pbar lithium lens was tripping on occasional ground fault.

Operations put Pbar in standby at 6:02 PM and put NuMI into a NuMI beam only mode.

At 6:09 PM, Operations reported that ARF1 and the sweeper were also tripping off.

At 6:17 PM, Operations reported that the I- Source extractor was sparking. Experts tuned on the source and it resumed normal operations.

A Pbar expert reported at 6:19 PM that there was a ground fault in the lens’ transformer.

At 6:39 PM, LRF2 tripped off. It reset.

Operations established store 5437 at 7:06 PM with an initial luminosity of 174.92E30.

A power supply expert reported at 7:13 PM that the lithium lens power supply was okay.

At Pbar expert began a stacktail measurement study at 8:32 PM with the 8mA of antiprotons in the stack.

At 10:52 PM, the TeV store 5437 appeared to have aborted due to a shunt and a fast bypass at TeV sector A2. There was no quench. (Later, it was determined that the initial cause due to the C2 TeV power supply.)

The Pbar study ended at 11:23 PM.

Operations reported at midnight that the Recycler had stashed once during the evening shift.

Saturday May 19

The midnight shift began with TeV experts investigating a power supply problem, with Operations monitoring the stash (47.0E10), with Pbar experts replacing the lithium lens unit, and with NuMI and MiniBooNE operating normally.

At 12:21 AM, LRF3 tripped off. It reset.

A TeV power supply expert began looking at the TeV sector C2 power supply at 1:44 AM.

At 2:52 AM, Operators tried to turn on the TeV, but the sector D3 VCB would not close. Operators reset the device and the VCB closed at 3:12 AM.

The Pbar lens work ended at 5:20 AM.

Pbar resumed stacking at 5:49 AM. Experts tuned.

Operations prepared MI-12B at 6:29 AM for and an access to check on a MiniBooNE dehumidifier water holding tank. Operations turned off MiniBooNE.

Operations reported at 8:18 AM that they accessed MI-12B and replaced a sump pump fuse. This didn’t fix the problem. Operators contacted MiniBooNE experts.

At 9:28 AM, the TeV suffered a sector A1 quench during TeV studies. Operators said that the cause of the quench was due to the Tevatron Electron Lens (TEL2) not being on; there was too much beam in the abort gap.

MiniBooNE experts reported at 11:27 AM that the pump for the holding tank was bad and that they were looking for a spare.

Operations reported at 12:34 PM that the TeV A1 cryo system was cooling down slowly due to poor vacuum.

Operations reported at 12:49 PM that the MiniBooNE de-watering pump had been replaced.

At 12:55 PM, Operations reported that the readback for the lithium lens cooling water system was broken. Operators called in an expert.

MiniBooNE resumed taking beam at 1:27 PM.

At 1:55 PM, Operations reported that they had completed a stash and the Pbar efficiencies were down. Operations informed experts.

Operations reported at 4:13 PM that the TeV A1 vacuum had recovered and the cryo system was cold.

A Pbar expert reported at 4:27 PM that he made changes to improve the transfer efficiency of the beam going from the Accumulator to MI.

Per request, Operators and the Duty Mechanic tried to turn on the TeV sector F1 pond pump at 5:28 PM. It wouldn’t start. Operations will notify the necessary people on Monday.

Operations put the TeV through a dry squeeze at 5:41 PM.

Operations put the TeV through a 4x0 wet squeeze at 6:39 PM.

Operations put the TeV into shot setup at 7:11 PM.

Operations reported at 8:09 PM that the Recycler had stashed four times during the last 12 hours. Accelerator status: The Pbar stack was at 27.0E10.

Operations established store 5444 at 8:59 PM with an initial luminosity of 149.7E30.

Sunday May 20

The midnight shift began with the TeV in shot setup, with Operations monitoring the stash (88.0E10), with Pbar stacking, and with NuMI and MiniBooNE taking beam.

At 1:15 AM, Operations reported that a Linac quadrupole power supply (L:Q43) was having regulation problems. Operations adjusted its voltage and this fixed the problem. This scenario occurred many more times over the rest of the shift.

At 6:25 AM, store 5444 aborted. There was no quench. Operations discovered an excessive voltage ripple from the C2 power supply. TeV experts were called in to investigate.

Operations reported at 8:04 AM that the Recycler had stashed five times during the last 12 hours. Accelerator status: the stack was at 24.9E10 and the stash was at 165.3E10.

A TeV expert reported at 10:46 AM that the C2 power supply might have two problems and one of them was a bad filter power supply. Operations called in an expert. The investigation continued.

Operations reported at 12:52 PM

that the TeV experts replaced the C2 voltage monitor chassis along with a cable. The TeV experts decided to run without the C2 power supply.

Operations put the TeV through the first of three ramps at 12:54 PM.

Around 1:10 PM, Operations reported that there was no Linac beam and no alarms. Operators turned off the Linac beam and investigated, but could not find anything wrong. When they turned the beam back on everything was okay.

Operations put the TeV through a dry squeeze at 2:12 PM.

At 3:02 PM, the TeV aborted at D2. Operators reported that a CAMAC crate at TeV sector D6 had failed.

At 3:04 PM, Operations reported that the Pelletron VME crate wasn’t responding. A reboot didn’t work; the Pelletron will need to be accessed.

Operators accessed the Pelletron at 3:59 PM and tried to reset the VME with no success. An expert came in to investigate.

Operations reported at 4 PM that they had replaced the failed D6 CAMAC crate.

Operations put the TeV through a 4x0 wet squeeze at 5:38 PM.

At 6:13 PM, the Linac beam went away again. This time it seemed to be due to the water resister column. Operators called in experts.

Operations put in an energy conservation time line into the Time Line Generator (TLG) at 6:41 PM.

Operations reported at 7:33 PM that controls experts had accessed and successfully rebooted the Pelletron VME.

Operations reported at 8:02 PM that the Recycler had stashed three times during the last 12 hours. Accelerator status: the stack was at 54.2E10 and the stash was at 267.4E10.

A Linac expert reported that they had found and fixed a bad resister that worked in parallel with the I- Source column’s water resister. Stacking and beam to NuMI resumed at 8:29 PM.

Operations put the TeV into shot setup at 8:30 PM.

After another round of L:Q43 problems, operators replaced the regulation card at 10:17 PM. This didn’t work. Operators changed the timing card and achieved stable running.

Operations established store 5450 at 11:22 PM with an initial luminosity of 165.7E30. This store had started colliding earlier, but D0 had a problem with beam positions so they held off turning up their voltages until those problems were resolved.

Monday May 21

The midnight shift began with Operations monitoring store 5450 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

At 12:15 AM, LRF4 tripped off. It reset.

Operations reported at 8:04 AM that the Recycler had stashed four times during the last 12 hours.

The Plans for Monday

The plans for today are to stack, stash, and store. The Run Coordinator may allow studies that have little impact on stacking and stashing.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 05/21/2007    email Fermilab